What is the angular velocity of a spinning object if it has a rotational kinetic energy of 280J, and a rotational inertia of 22900kg*m2?
Answer #1

The rotational Kinetic energy is given as

K.E = 1/2 * I * w2

where I is Inertia and w is angular velocity

280 = 1/2 * 22900 * w2

w2 = 2 * 280 / 22900

w2 = 0.0244

so,

w = 0.15637 rad/sec